what is the final delivered electricity in kWh by burning one pound of coal (13,100 BTU) with a boiler loss of 20%, generator loss of 50%, and transmission loss of 2%?
1
Expert's answer
2012-09-07T11:00:00-0400
Q0 = 13,100 BTU = 13,821,231.7 J.
1 kWh = 3.6 MJ, so Q0 = 3.839 kWh.
The
final delivered electricity Qf = Q0 * (100 - Lb - Lg - Lt) / 100 =
3.839 *
(100 - 20 - 50 - 2) / 100 = 1.075 kWh.